1) Tom the Turnip Man is having trouble deciding how to run his business and has...

1)

Tom the Turnip Man is having trouble deciding how to run his business and has asked me to ask you what he should do. Tom is a price-taker and receives $10 for every box of turnips he sells. If he operates at a point where price is equal to his marginal cost, he will sell 1000 boxes of turnips. Unfortunately his total cost of doing business at that output level is $15,000, of which $7,000 are total variable costs. What should Tom do in the short run?

Group of answer choices

a Stop producing turnips

b Raise the price of his turnips

c Sell more turnips at the same price

d Continue producing 1000 boxes of turnips

2)

In which of the following market structures will we always find advertising?

Group of answer choices

a Monopolistic Competition

b Monopoly

c Oligopoly

d Perfect Competition
Answer #1

1) d Continue producing 1000 boxes of turnips
(AVC = VC/Q = 7000/1000 = 7
So, P > AVC which means that firm should keep producing in the short run.)

2) c Oligopoly
(Advertising is highly concentrated in oligopoly markets.)